OpenStudy (v-sync):

If the area of a square is 1024 m2, find the perimeter.

OpenStudy (v-sync):

1024/4 = 256 265 * 4 = 1024 area = perimeter on squares? @Will.H

OpenStudy (will.h):

\[A = a^2\] --> are of a square we will need to find the length of the side we may do that through transfering the formula above to \[a = \sqrt A\] where A is the area

OpenStudy (will.h):

So how much would the length of side be?

OpenStudy (will.h):

need help?

OpenStudy (v-sync):



OpenStudy (will.h):

well you are wrong solving that way the length of the side is not 4! solve using the formula i posted for solving for the side \[a = \sqrt A\] where A is the area so therefor \[a = \sqrt 1024\] which sets a = 32 <--- that's the length of the side Now here's what we have 

OpenStudy (will.h):

Now you can find the perimeter

OpenStudy (v-sync):

I know the length of th eside is not 4, that was an example

OpenStudy (will.h):

oh then your example is correct but not always the piremeter and area of a square are the same

OpenStudy (will.h):

in fact it is rare

OpenStudy (will.h):

let me know if you need additional help

OpenStudy (v-sync):

√1024 = 32 32 * 4 = 128 perimeter = 128

OpenStudy (will.h):

Indeed